# Laypeople’s Egocentric Perceptions of Copyright for AI-Generated Art

### v1

@@ 4.2.1. User
When judging others’ submissions, invested evaluators attributed more authorship to creators than creators attributed to themselves (diff = ). This effect is the opposite of the one hypothesized in H1. Moreover, invested evaluators attribute more authorship to others than to themselves (diff = ). This effect goes against our hypothesis H2.
We also observed some order effects in participants’ attribution of authorship to the user. The more images a participant evaluated, the more authorship they attributed to users ( = 0.050, = 0.021, = , ¡ .05). We note that the results described above are robust to including the order variable as a covariate. Invested evaluators attributed more authorship to creators (i.e., users) than creators themselves (diff = ). Furthermore, invested evaluators attributed more authorship to others than to themselves (diff = ).
@@ 4.2.2. Other Entities
There were no significant differences in the perceived authorship of the AI model, the developer, and data contributors across treatments. Nonetheless, we note a borderline significant order effect on perceived authorship of the AI model, such that the more images participants evaluated, the less authorship they attributed to the AI model ( = -0.056, = 0.030, = , = 0.061).
@@ 4.3. RQ3: Attribution of Rights
Figure 5 presents participants’ opinions regarding who should hold the rights to display and make copies of AI-generated images. Users (i.e., those who used the GenAI to generate images) were selected by more than 60% of participants for all rights, in both commercial and non-commercial settings. Data contributors were identified as rights-holders by approximately 50% of participants across all rights and settings. The company that developed the AI model was granted the rights to display and make copies of AI-generated images by around 41% of participants in non-commercial settings; however, only 32% of respondents believed the company should have commercial rights over the image. The AI model was recognized as a rights-holder by around 19% and 27% of participants for commercial and non-commercial uses, respectively.
Participants were also able to indicate that no one should have exclusive rights over the images by selecting that “anyone” should be able to display and make copies of AI-generated images. We observed clear differences in responses between commercial and non-commercial rights. Participants were more likely to support non-exclusive rights in non-commercial settings (approximately 57% and 47% in favor of anyone having the right to display and make copies, respectively). In contrast, when evaluating commercial rights, only a few respondents indicated that anyone should have them (approximately 9%).
Unlike for attributions of authorship, our treatments had little effect on attributions of rights. That is, the patterns described above are fairly consistent across all treatments. Hence, for brevity, we omit the pairwise comparisons tables from the paper and comment on all of the differences identified as statistically significant directly in the text below.
We did not find support for any of our hypotheses regarding attribution of rights to users. There were only a few significant pairwise differences between conditions. Creators were relatively more likely to support non-exclusive rights to display their own creations than invested evaluators evaluating the same images (OR = 4.347, = 3.532, p ¡ .005 for commercial use, OR = 2.253, = 3.043, p ¡ .05 for non-commercial use). When evaluating other people’s creations, invested evaluators were less likely to support non-exclusive rights to display images commercially than uninvested evaluators (OR = 0.340, = -2.801, p ¡ .05). Finally, creators were borderline more supportive of granting the right to make copies commercially to data contributors than uninvested evaluators (OR = 0.497, = -2.622, p ¡ 0.05). No other differences were statistically significant.
@@ 4.4. Score Evaluation
| Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1515) = 8.686 | |||
| Creators (Own) - Invested Evaluators (Others) - H1 | t(1515) = 8.425 | |||
| Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1515) = 8.264 | |||
| Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1515) = -0.924 | |||
| Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1515) = 7.807 |
Figure 6 presents participants’ mean score evaluations on a 11-point scale (0 = Very bad, 10 = Very good). These scores were used to determine who was granted the monetary award and which images were displayed in our AI art exhibition. On average, participants evaluated images slightly positively (mean = 6.61, = 2.38). Table 3 presents pairwise comparisons between treatment conditions.
Pairwise comparisons between treatments indicate clear egocentric effects. In line with H1, creators evaluated their own creations more highly than invested and uninvested evaluators judging the same images (diff = and diff = , respectively). As hypothesized by H2, invested evaluators judged their own images more positively than those created by other participants (diff = ). We found that uninvested evaluators score images generated by creators lower than invested evaluators score their own creations (diff = ), as suggested by H4.
@@ 5.1.1. RQ1: Creativity and Effort—Not Skills—Are Required To Generate Art with AI
Participants believed that creativity and effort were necessary to generate images with the GenAI model. In contrast, they neither agreed nor disagreed that creators used their skills when creating AI art.
The fact that the AI generation process was described as creative seems to suggest that AI art could have the modicum of creativity that US copyright law requires for a work to warrant copyright (Office, 2023), which is at odds with current decisions to reject protection for AI-generated images (Davis, 2023). Concerning judgments of effort, we note that our study provided evaluators with information that can work as a proxy for effort (e.g., the total number of images creators generated and how long they spent in the study), which could have influenced how participants judged the effort put by creators. Nonetheless, our experimental findings are robust to the inclusion of these covariates. Finally, perceptions of skills could have been limited by how much control creators had over the images they generated. Had creators been allowed to edit the GenAI outputs, evaluators could have judged the creation process as requiring more skills—a research direction that future work could explore.
@@ 5.1.2. RQ2: Users and Data Contributors as Authors
Users and data contributors were attributed the most authorship for AI-generated art. Although it is less surprising that those who prompted the GenAI model to generate images are perceived as authors, participants’ acknowledgement of data contributors is noteworthy. In line with prior work suggesting that laypeople agree that data contributors should be compensated if their work is used for training GenAI models (Kastrenakes and Vincent, 2023), our results suggest that current practices that fail to compensate and appreciate data contributors, which some would equate to theft (Chayka, 2023; Merkley, 2023), may be at odds with laypeople’s expectations. People’s opinions seem to be more aligned with proposals of licensing models, under which data contributors can be compensated for the use of their creations (Butler, 2024).
@@
Our result that users were acknowledged as authors of the images they generate using AI calls for the reconsideration of legal decisions that have refused to grant copyright to GenAI ouputs in the US (Davis, 2023). The main reason given for the rejection was that AI-generated images do not have human authors, making it ineligible for copyright protection. In contrast, our findings suggest that laypeople believe users are authors, even in a scenario in which they were not allowed to edit the AI-generated images. We expect the perceived authorship of users to increase when given the chance to further exert control over GenAI outputs.
Participants were uncertain about the authorship of the AI model, which is aligned with prior work suggesting that laypeople are neither against nor in favor of copyright rights to AI models (Lima et al., 2020). Finally, respondents’ disagreement with the idea of companies being authors of GenAI outputs are at odds with potential interpretations of UK IP law, which could be interpreted as the companies training GenAI models holding copyright over outputs (Ihalainen, 2018). That is, current practices that concentrate profits in the hands of the corporations training GenAI models without little to no compensation to data contributors (Chayka, 2023; Merkley, 2023) do not align with lay opinions about copyright of AI-generated art.
We note, however, these these two actors, to whom participants attributed lower levels of authorship, are non-human entities. In contrast, the user and data contributors are human, suggesting that our findings could have been influenced by the human nature of these actors. Future work could replicate our study by also investigating whether the (human) programmers that developed the model would be granted authorship, which could indicate that the human nature of potential authors plays a role in people’s perceptions of authorship.
@@ 5.1.3. RQ3: Users and Data Contributors as Rights-Holders
Similar to our results concerning authorship, those who prompted the AI model to generate images (i.e., users) and data contributors were frequently identified as potential rights-holders of AI-generated art. This finding has two main implications. First, it puts into question the exclusive nature of the rights afforded by copyright protection; participants seemed to support a more distributed approach, under which not only “authors” would be owners of the copyright, but also some actors without which AI-generated art would not be possible. For instance, a joint ownership model (Khosrowi et al., 2023) could be more aligned with laypeople’s expectations regarding AI-generated art. Second, it highlights participants’ calls for the compensation of data contributors. An approach that could be more aligned with lay opinions could rely on neighboring rights (Lee, 2021; Iaia, 2022), which would still center ownership around “authors” (i.e., users) but without neglecting the interests of data contributors, e.g., through the distribution of royalties. As discussed above, another solution could be based on licensing models for training data (Butler, 2024).
Participants were more supportive of non-exclusive rights over AI-generated art in non-commercial settings, in line with doctrines that permit the use of copyrighted material for non-commercial purposes (e.g., the US fair use doctrine). Future debates concerning the use of AI-generated outputs in relation to copyright could consider our results to determine whether their use should be legal.
@@ 5.2. Egocentric Biases (Or Lack Thereof) in Perceptions of AI-Generated Images
5.2.1. Creativity, Effort and Skills
Judgments regarding creativity and effort partially exhibited egocentric effects. Creators attributed more creativity to their creations than uninvested evaluators (partial H1), in line with prior work (Buccafusco and Sprigman, 2011). Evaluations of the effort put in generating images with GenAI are also consistent with our egocentric hypotheses (partial H1 and H2). Furthermore, judgments concerning creativity are aligned with H4 (i.e., experience effect), which suggests that people may judge images more favorably after interacting with the GenAI model. Judgments about effort and skills are also partially aligned with H4.
@@ 5.2.2. Perceived Authorship
A potential explanation to why creators attributed less authorship to themselves than to other people is that authorship does not only encode positive rights but also negative responsibilities from which creators tried to distance themselves. Although authorship is often associated with positive outcomes in the context of copyright law (e.g., the opportunity to profit from one’s creation), it could also lead to the responsibility for any negative outcomes or backlash arising from their work being published (Nehamas, 1986; Foucault, 2019). If one is an author under this interpretation, they are also responsible, for instance, if their creation is found to be infringing on another copyrighted work, which is relevant for debates concerning AI-generated images. Hence, creators seem to want to enjoy the rights associated with copyright (see above), but might rather not be perceived as authors because of potential negative repercussions in case something goes wrong. Future work could explore this potential explanation further.
@@
In H4, we hypothesized that having experience with the AI model would influence participants’ attribution of authorship to the user. Informally, we expected that by interacting with the AI model, people would notice the degree of influence a user’s prompts have over the model’s outputs, leading them to attribute more authorship to users. We did not, however, find a significant difference between uninvested evaluators and other groups of participants for perceptions of user authorship. It is important to note that H4 focused on the experience of observing the relationship between one’s own inputs and the model’s outputs. Perhaps there was not much variance in their own inputs, or they generated too few images (the average participant prompted the model a median of two times) for the experience effects to kick in.
@@
However, we identified a different effect of experience. The more images a participant evaluated, the more authorship they attributed to users, suggesting order effects. The identified order effects relate to a different form of experience: observing the relationship between other participants’ inputs and model outputs. It is likely that there was more between-participants variance than within-participant variations in the images they generated.666This would be aligned with prior research in social psychology, which finds higher degrees of variance in the judgments of different people than between the judgments of the same person in different points in time (Kahneman et al., 2016, 2021) Since evaluators observed the prompts from four different creators, they had more opportunities for the effect of experience to come into play.
@@
A potential explanation connecting our findings regarding H1, H2, H4 and order effects is that being exposed to diverse GenAI inputs and outputs makes evaluators value the human component of AI-generated images more. After using the GenAI model, seeing what other participants generated, and realizing that different human actions led to widely different outputs, invested evaluators may have started acknowledging that AI cannot generate images without some level of human input and authorship. Creators and invested evaluators judged their own images without such experience, while invested evaluators judged the creators’ images after this experience, hence possibly explaining why the latter received higher authorship scores (opposite of H1 and H2). To explore this potential explanation, future work could explore inverting the order in which invested evaluators generate and evaluate images. By first evaluating other people’s submissions and then generating their own image, participants might end up overestimating their authorship after seeing what others were able to create, providing further evidence of order effects.
@@
Finally, we note that the increase in the perceived authorship of the user was accompanied with a slight decrease in AI model’s authorship. A potential factor influencing this shift in authorship could be the perceived skills of the user, which exhibit a similar pattern across our experimental conditions. It is possible that seeing how different participants generated images of varying quality—while using the same GenAI model—highlighted the skills behind prompting, leading to higher perceived authorship to users and a slightly lower perceived authorship of the AI model.

### v2

@@ 4.2.1. User
When judging others’ submissions, invested evaluators attributed more authorship to creators than creators attributed to themselves (diff = ). This effect is the opposite of the one hypothesized in H1. Moreover, invested evaluators attributed more authorship to others than to themselves (diff = ). This effect goes against our hypothesis H2. Finally, we found that invested evaluators attributed more authorship to creators than uninvested evaluators (diff = ), an effect in the opposite direction to that hypothesized in H3 and partially in line with H4.
We also observed borderline significant order effects in participants’ attribution of authorship to the user. The more images a participant evaluated, the more authorship they attributed to users ( = 0.041, = 0.022, = , = .058). We note that the results described above are robust to including the order variable as a covariate. Invested evaluators attributed more authorship to creators (i.e., users) than creators themselves (diff = ). Invested evaluators attributed more authorship to others than to themselves (diff = ). However, the statistically significant difference in how invested and uninvested evaluators attributed authorship to creators disappeared when accounting for order effects ( = 0.068).
@@ 4.2.2. Other Entities
There were no significant differences in the perceived authorship of the AI model, the developer, and data contributors across treatments. Nonetheless, we note a significant order effect on perceived authorship of the AI model, such that the more images participants evaluated, the less authorship they attributed to the AI model ( = -0.064, = 0.028, = , ¡ 0.05).
@@ 4.3. RQ3: Attribution of Rights
The figure consists of four panels. Each panel shows bar plots that represent the percentage of participants who picked a particular actor as the holder of a right. Actors are represented in the x-axis. The top-left panel corresponds to the right to display images commercially, the top-right panel refers to the right to display non-commercially, the bottom-left panel corresponds to the right to make copies commercially, and the bottom-right refers to the right to make copies non-commercially. Each panel also shows the data separated by experimental condition.
Figure 5 presents participants’ opinions regarding who should hold the rights to display and make copies of AI-generated images. Users (i.e., those who used the GenAI to generate images) were selected by more than 60% of participants for all rights, in both commercial and non-commercial settings. Data contributors were identified as rights-holders by approximately 50% of participants across all rights and settings. The company that developed the AI model was granted the rights to display and make copies of AI-generated images by around 41% of participants in non-commercial settings; however, only 33% of respondents believed the company should have commercial rights over the image. The AI model was recognized as a rights-holder by around 19% and 27% of participants for commercial and non-commercial uses, respectively.
Participants were also able to indicate that no one should have exclusive rights over the images by selecting that “anyone” should be able to display and make copies of AI-generated images. We observed clear differences in responses between commercial and non-commercial rights. Participants were more likely to support non-exclusive rights in non-commercial settings (approximately 58% and 47% in favor of anyone having the right to display and make copies, respectively). In contrast, when evaluating commercial rights, only a few respondents indicated that anyone should have them (approximately 9%).
Unlike for attributions of authorship, our treatments had little effect on attributions of rights. That is, the patterns described above are consistent across all treatments, with a few exceptions detailed below. Hence, for brevity, we report the full pairwise comparison tables in the Appendix and report all of the differences identified as statistically significant directly in the text below.
We did not find support for any of our hypotheses regarding attribution of rights to users. There were only a few significant pairwise differences between conditions. Creators were relatively more likely to support non-exclusive rights to display their own creations than invested evaluators evaluating the same images (OR = 3.953, = 3.330, p ¡ .005 for commercial use, OR = 2.193, = 2.941, p ¡ .05 for non-commercial use). Finally, creators were borderline less supportive of granting the right to make copies commercially to data contributors than uninvested evaluators (OR = 0.502, = -2.589, p ¡ 0.05). No other differences were statistically significant.
@@ 4.4. Score Evaluation
The figure depicts a bar plot that illustrates the mean score evaluation across all images in our study. The plot also shows mean values depending on the treatment condition.
@@
Creators (Own) - Uninvested Evaluators (Others) - H1 and H4 | t(1534) = 8.821 | |||
Creators (Own) - Invested Evaluators (Others) - H1 | t(1534) = 8.189 | |||
Invested Evaluators (Own) - Invested Evaluators (Others) - H2 | t(1534) = 8.232 | |||
Uninvested Evaluators (Others) - Invested Evaluators (Others) - H3 and H4 | t(1534) = -1.242 | |||
Invested Evaluators (Own) - Uninvested Evaluators (Others) - H4 | t(1534) = 8.037 |
Figure 6 presents participants’ mean score evaluations on an 11-point scale (0 = very bad, 10 = very good). These scores were used to determine who was granted the monetary award and which images were displayed in our AI art exhibition. On average, participants evaluated images slightly positively (mean = 6.62, = 2.40). Table 3 presents pairwise comparisons between treatment conditions.
Pairwise comparisons between treatments indicate clear egocentric effects. In line with H1, creators evaluated their own creations more highly than invested and uninvested evaluators judging the same images (diff = and diff = , respectively). As hypothesized by H2, invested evaluators judged their own images more positively than those created by other participants (diff = ). We found that uninvested evaluators score images generated by creators lower than invested evaluators score their own creations (diff = ), as suggested by H4. Finally, we also observed significant order effects. The more images a participant evaluated, the lower they scored images ( = -0.094, = 0.046, = , ¡ 0.05).
@@ 5.1.1. RQ1: Creativity and Effort—Not Skills—Are Required To Generate Art with AI
Participants believed that creativity and effort were necessary to generate images with the GenAI model. In contrast, they neither agreed nor disagreed that creators used their skills when creating AI art. In other words, when we analyze how participants perceived the nature of the digital artifacts produced by GenAI through the lens of these copyright-relevant factors, we find that people believe AI-generated art reproduces its creator’s creativity and effort, even though producing it may not require advanced artistic skills.
The fact that the AI generation process was described as creative seems to suggest that AI art could have the modicum of creativity that US copyright law requires for a work to warrant copyright (Office, 2023), which is at odds with current decisions to reject protection for AI-generated images (Davis, 2023). Concerning judgments of effort, we note that our study provided evaluators with information that can work as a proxy for effort (e.g., the total number of images creators generated and how long they spent in the study), which could have influenced how participants judged the effort put by creators. Nonetheless, our experimental findings are robust to the inclusion of these covariates. That is, our findings remain qualitatively the same even when accounting for the aforementioned proxies for effort. Finally, perceptions of skills could have been limited by how much control creators had over the images they generated. Had creators been allowed to edit the GenAI outputs, evaluators could have judged the creation process as requiring more skills—a research direction that future work could explore.
@@ 5.1.2. RQ2: Users and Data Contributors as Authors
Users and data contributors were attributed the most authorship for AI-generated art. Although it is less surprising that those who prompted the GenAI model to generate images are perceived as authors, participants’ acknowledgment of data contributors is noteworthy. In line with prior work suggesting that laypeople agree that data contributors should be compensated if their work is used for training GenAI models (Kastrenakes and Vincent, 2023), our results suggest that current practices that fail to compensate and appreciate data contributors—which some equate to theft (Chayka, 2023; Merkley, 2023)—may be at odds with laypeople’s expectations. People’s opinions seem to be more aligned with proposals of licensing models, under which data contributors can be compensated for the use of their creations (Butler, 2024).
@@
By designing systems that fail to compensate those whose creations are needed to train GenAI, system designers may end up perpetuating existing patterns of—what some authors would call—exploitation (Chayka, 2023; Merkley, 2023). Worryingly, past work suggests that even when developers want to question the implications of their products, they often lack the power to contest organizational dynamics and incentives (Widder et al., 2023, 2024). Future work could bring together prior research developing mechanisms for collective action (Krafft et al., 2021; Calacci and Pentland, 2022) together with self-questioning design practice and education on the potential ethical concerns posed by GenAI (Garrett et al., 2020) to identify mechanisms through which HCI practitioners and researchers can resist entrenchment of exploitive design practices.
@@
Our result that users were acknowledged as authors of the images they generate using AI calls for the reconsideration of legal decisions that have refused to grant copyright to GenAI outputs in the US (Davis, 2023). The main reason given for the rejection was that AI-generated images do not have human authors, making them ineligible for copyright protection. In contrast, our findings suggest that laypeople believe users are authors, even in a scenario in which they were not allowed to edit the AI-generated images. We expect the perceived authorship of users to increase when given the chance to exert more control over GenAI outputs.
Participants were uncertain about the authorship of the AI model, which is aligned with prior work suggesting that laypeople are neither against nor in favor of copyright rights to AI models (Lima et al., 2020). Finally, respondents’ disagreement with the idea of companies being authors of GenAI outputs is at odds with potential interpretations of UK IP law, which could be interpreted as the companies training GenAI models holding copyright over outputs (Ihalainen, 2018). That is, current practices that concentrate profits in the hands of the corporations training GenAI models without little to no compensation to data contributors (Chayka, 2023; Merkley, 2023) do not align with lay opinions about copyright of AI-generated art.
We note, however, that these two actors, to whom participants attributed lower levels of authorship, are non-human entities. In contrast, the user and data contributors are human, suggesting that our findings could have been influenced by the human nature of these actors. Future work could replicate our study by also investigating whether the (human) programmers who developed the model would be granted authorship, which could indicate that the human nature of potential authors plays a role in people’s perceptions of authorship.
@@ 5.1.3. RQ3: Users and Data Contributors as Rights-Holders
Similar to our results concerning authorship, those who prompted the AI model to generate images (i.e., users) and data contributors were frequently identified as potential rights-holders of AI-generated art. All in all, we find that people grant ownership of AI-generated digital artifacts to users and those who contributed creatively to the system design and not necessarily to the technological artifact itself and the entity who developed it.
Our findings have two main legal implications. First, it puts into question the exclusive nature of the rights afforded by copyright protection; participants seemed to support a more distributed approach, under which not only “authors” would be owners of the copyright, but also some actors without which AI-generated art would not be possible. For instance, a joint ownership model (Khosrowi et al., 2023) could be more aligned with laypeople’s expectations regarding AI-generated art. Second, it highlights participants’ calls for the compensation of data contributors. An approach that could be more aligned with lay opinions could rely on neighboring rights (Lee, 2021; Iaia, 2022), which would still center ownership around “authors” (i.e., users) but without neglecting the interests of data contributors, e.g., through the distribution of royalties. As discussed above, another solution could be based on licensing models for training data (Butler, 2024).
Participants were more supportive of non-exclusive rights over AI-generated art in non-commercial settings, which is in line with doctrines that permit the use of copyrighted material for non-commercial purposes (e.g., the US fair use doctrine). Future debates concerning the use of AI-generated outputs in relation to copyright could consider our results to determine whether their use should be legal.
Although copyright law worldwide does not yet have a definitive answer to whether AI-generated art is copyrightable, another legally relevant document that might determine who might own outputs is each GenAI model’s terms of service. Even though most terms of service state that users own the content they create “to the extent permitted by applicable law” (OpenAI, 2024; Midjourney, 2024; Stability.AI, 2022), they also grant companies the right to use the AI-generated content to improve their models. Such terms could acknowledge the role of data contributors, restrict the rights of developing companies, and consider whether the content is being used commercially or not.666Midjourney (Midjourney, 2024), for instance, only grants ownership of AI outputs to employees of companies with more than $1,000,000 USD if they are subscribed to one of their premium plans.
@@ 5.2. Egocentric Biases (Or Lack Thereof) in Perceptions of AI-Generated Images
5.2.1. Creativity, Effort, and Skills
Judgments regarding creativity and effort partially exhibited egocentric effects. Creators attributed more creativity to their creations than uninvested evaluators (partial H1), in line with prior work (Buccafusco and Sprigman, 2011). Evaluations of the effort put into generating images with GenAI are also consistent with our egocentric hypotheses (partial H1 and H2). Furthermore, judgments concerning creativity are aligned with H4 (i.e., experience effect), which suggests that people may judge images more favorably after interacting with the GenAI model. Judgments about effort and skills are also partially aligned with H4.
@@ 5.2.2. Perceived Authorship
A potential explanation for why creators attributed less authorship to themselves than to other people is that authorship does not only encode positive rights but also negative responsibilities from which creators tried to distance themselves. Although authorship is often associated with positive outcomes in the context of copyright law (e.g., the opportunity to profit from one’s creation), it could also lead to the responsibility for any negative outcomes or backlash arising from their work being published (Nehamas, 1986; Foucault, 2019). If one is an author under this interpretation, they are also responsible, for instance, if their creation is found to be infringing on another copyrighted work, which is relevant for debates concerning AI-generated images. Hence, creators seem to want to enjoy the rights associated with copyright (see above) but might rather not be perceived as authors because of potential negative repercussions in case something goes wrong. Future work could explore this potential explanation further. As suggested by prior work showing that users can escape responsibility by delegating tasks to AI systems (Feier et al., 2022; Köbis et al., 2021), studies could examine how creators distance themselves from AI-generated art when negative responsibilities are salient.
Participants’ distancing from potential negative outcomes of using GenAI is in conflict with current terms of service. Existing terms state that users are fully responsible for whatever they generate using the model, including if outputs infringe on the copyright of any third party (OpenAI, 2024; Stability.AI, 2022; Midjourney, 2024). Companies developing GenAI models could consider decreasing (or sharing) the potential liability of users in their terms of service to align with users’ expectations.
@@
In H4, we hypothesized that having experience with the AI model would influence participants’ attribution of authorship to the user. Informally, we expected that by interacting with the AI model, people would notice the degree of influence a user’s prompts have over the model’s outputs, leading them to attribute more authorship to users. We did not, however, find a significant difference between uninvested evaluators and other groups of participants in terms of perceptions of user authorship. It is important to note that H4 focused on the experience of observing the relationship between one’s own inputs and the model’s outputs. Perhaps there was not much variance in each participant’s inputs, or they generated too few images (the average participant prompted the model a median of two times) for the experience effects to kick in.
@@
However, we identified a different effect of experience. The more images a participant evaluated, the more authorship they attributed to users, suggesting order effects. The identified order effects relate to a different form of experience: observing the relationship between other participants’ inputs and model outputs. It is likely that there was more between-participants variance than within-participant variations in the images they generated.777This explanation would be aligned with prior research in social psychology, which finds higher degrees of variance in the judgments of different people than between the judgments of the same person in different points in time (Kahneman et al., 2016, 2021). Since evaluators observed the prompts from four different creators, they had more opportunities for this effect of experience to come into play.
@@
A potential explanation connecting our findings regarding H1, H2, H4, and order effects is that being exposed to diverse GenAI inputs and outputs makes evaluators value the human component of AI-generated images more. After using the GenAI model, seeing what other participants generated, and realizing that different human actions led to widely different outputs, invested evaluators may have started acknowledging that AI cannot generate images without some level of human input and authorship. Creators and invested evaluators judged their own images without such experience, while invested evaluators judged the creators’ images after this experience, hence possibly explaining why the latter received higher authorship scores (opposite of H1 and H2). To explore this potential explanation, future work could explore inverting the order in which invested evaluators generate and evaluate images. By first evaluating other people’s submissions and then generating their own image, participants might end up overestimating their authorship after seeing what others were able to create, providing further evidence of order effects.
@@
Finally, we note that the increase in the perceived authorship of the user was accompanied by a slight decrease in AI model’s authorship. It is possible that seeing how different participants generated images of varying quality—while using the same GenAI model—highlighted the skills behind prompting, leading to higher perceived authorship to users and slightly lower perceived authorship of the AI model.